equivalent weight of pottash alum is M1/x and euivalent weight of gypsum is M2/y .then x-y will be ?

Grade:12

1 Answers

Vikas TU
3 years ago
For question we have x-y=6
●We know that potash alum formula is K2SO4.Al2(SO4)3.24H20
The molar mass of potash alum is 948.73 grams and for potash alum n factor is 8 as 2 potassium 4 sulphate ion and 2 aluminium ions so total ions are 8 hence n factor is 8.
Equivalent weight=molecular weight/nfactor=M1/x
So x=8
●Now for gypsum formula CaSO4.2H2O the n factor is 2 as 1 calcium and 1 sulphate ion so
Equivalent weight=molecular weight/n factor=M2/y
So y =2
●Now x-y=8-2=6